OpenStudy (anonymous):

Which arrangement for these numbers is from least to greatest? -5/8, -7/9, -4/5, -3/7 A.-7/9, -5/8, -3/7, -4/5 B. -4/5, -7/9, -5/8, -3/7 c.-4/5, -5/8, -7/9, -3/7 d. -3/7, -4/5, -5/8, -7/9

OpenStudy (welshfella):

i find these easier if i convert all the fractions to decimal use your calculator so for example -7/9= -0.7778

OpenStudy (anonymous):

Do i just divide the top and bottom numbers?

OpenStudy (anonymous):

4/7 I think.

OpenStudy (anonymous):

My possibilities were 6/7, 4/5, 4/7, & 11/6

OpenStudy (anonymous):

i got it wrong, it was 6/7 :/

OpenStudy (welshfella):

divide yop by bottom

OpenStudy (welshfella):

-6/7 = -0.857 -7/9 = -0.778 4/5 = -0.8 -3/7 = -0.429

OpenStudy (welshfella):

oh ne -6/7 -5/8 = -0.625

OpenStudy (welshfella):

-4/5 is the lowest

OpenStudy (welshfella):

B is the correct option

OpenStudy (welshfella):

remember we are dealing with negative numbers so -0.8 is less thnn -0.7